Solve the system of equations by using the addition method
x + 7y = 69
-2x + 7y = 51
Please show steps.
Thank you.

Answer by Alan3354(69443)

Solve the system of equations by using the addition method
x + 7y = 69
-2x + 7y = 51
------------------ Subtract